# A note for Carleson measure on bounded $\mathbb{C}$}-doconvex domains

### v1

@@ Abstract.
Let $0<q<p<\infty$, $\Omega$ be a bounded $\mathbb{C}$}-convex domains in ${\mathbb{C}}^{n}$. We establish several equivalent characterizations for the boundedness of Carleson embedding $J_{\mu}:A_{\alpha}^{p}\hookrightarrow L^{q}(\mu)$ on $\Omega$ with sharp $\mathcal{B}$} condition. Furthermore, we prove that the boundedness of $J_{\mu}$ is equivalent to its compactness.
@@ 1. Introduction and main result
The extension of $q$-Carleson measure to Bergman spaces has attracted considerable attention from many scholars, it was characterized by Hastings [17] and Oleinik [18] for $p\leq q$, Luecking [16] for $p>q$. In higher dimension, Cima-Wogen [7] and Luecking-Duren-Weir [14, 10] stuied the case of unit ball ${\mathbb{B}}^{n}$ by using the Euclidean geometry of ${\mathbb{B}}^{n}$ and conformal invariants (such as Bergman or Kobayashi balls) repectively. In order to extend the above results to other domains of ${\mathbb{C}}^{n}$, Cima-Mercer and Abate-Raissy-Saracco characterrized the $q$-Carleson measure for $A^{p}$ on strongly pseudoconvex domains by using the Euclidean geometric quantities and intrinsic Kobayashi geometry respectively, see [1, 2, 3, 8] for the details. For the case of weakly pseudoconvex domains, Li-Liu-Wang [13] characterized the $2$-Carleson measurable for $A^{2}$ by using Kobayashi balls on bounded convex domains with smooth boundary of finite type. Bi-Su-Zhang [5] characterized the $2$-Carleson measure for $A^{2}$ and $A^{p}$ respectively on bounded $\mathbb{C}$}-convex domains.
Khanh-Liu-Thuc [11] introduced the sharp $\mathcal{B}$} type condition (see Sect.2.4 ) for Bergman kerenl. Roughly speaking, Bergman kernel satisfies the sharp $\mathcal{B}$} condition ensures that the Bergman projection is bounded on $L^{p}$ for $1<p<\infty$. More recently, there is a series of work involved function spaces and operator theory related such domains. For example, Khanh-Liu-Thuc-Tien [11, 12] characterrized the boundedness and Schatten class of Bergman-Toeplitz operators for weighted spaces on weakly pseudoconvex domains with Bergman kernel satisfies sharp $\mathcal{B}$} condition. Zhang [19] characterized the $q$-Carleson measure for the weighted Bergman spaces $A_{\alpha}^{p}$ on bounded $\mathbb{C}$}-convex domains with sharp $\mathcal{B}$} type condition, and what’s more, Zhang’s results extended the results in [5] to the case of $p\leq q$ and the weighted Bergman spaces $A_{\alpha}^{p}$, but for the case of $p>q$, Zhang didn’t characterized the full Carleson measure, Zhang’s result is stated as follows.
@@
[19, Theorem 1.8] Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel of sharp $\mathcal{B}$} type, $\mu$ be a finite positive Borel measure on $\Omega$. Assume that $0<q<p<\infty$, if $\mu$ satisfies one of the following three conditions
@@
Let $\Omega$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel of sharp $\mathcal{B}$}-type. Let $\mu$ be a finite positive Borel measure, $0<q<p<\infty$. Then the following are equivalent.
@@ 2.2. Invariants
$\kappa_{\Omega}(x;v):=\inf\bigl\{|\xi|:f\in{\mathcal{H}}(\mathbb{D},\Omega),\text{ with }f(0)=x,\;(df)_{0}(\xi)=v\bigr\}.$
For any curve $\sigma:[a,b]\to\Omega$, the Kobayashi length is
@@
Suppose $\Omega$ is a bounded domains in ${\mathbb{C}}^{n}$, $z\in\Omega$ and $r\in(0,1)$. Let $B_{\Omega}(z,r)$ be the Kobayashi balls of centre $z$ and radius $\frac{1}{2}\log\frac{1+r}{1-r}$, i.e.
@@
For a bounded pseudoconvex domains $\Omega\subset{\mathbb{C}}^{n}$, Set $A^{2}$ be the Hilbert spaces consists of all $f\in{\mathcal{H}}(\Omega)$ such that $\|f\|_{A^{2}}^{2}=\int_{\Omega}|f(z)|^{2}dV(z)<\infty$, where $dV(z)$ is the ordinary Lebesgue measure. Riesz’s theorem yields that there exists a unique reproducing kernel function $K_{\Omega}(z,w)$ such that
@@ 2.3. Geometry of $\mathbb{C}$}-convex domains
and. Continuing this procedure we are led to an orthonormal basis
@@
[NIkolv2011, Theorem 1] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines, and $z\in\Omega$. Then for any $r\in(0,1)$,
@@
[5, Proposition 2.3] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. For any $r\in(0,1)$, we denote $R:=\frac{1}{2}(1+r)\in(0,1)$. Then for any $z_{0}\in\Omega$ and $z\in B_{\Omega}(z_{0},r)$, we have
@@
The following cover type lemma plays an significant role in the proof of Theorem 1.2.
@@
[19, Lemma 2.4] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. Then for any $r\in(0,1)$, $z\in\Omega$ and $w\in B_{\Omega}(z,r)$, we have that
@@
For $z\in\Omega$ and $r\in(0,1)$. Let $p_{\Omega}(z)=\prod\limits_{i=1}^{n}\eta_{i}(z)$, then by [19], we have
(2.1) $\frac{1}{C}p_{\Omega}(z)\leq p_{\Omega}(w)\leq Cp_{\Omega}(z),\quad w\in B_{\Omega}(z,r)$
for some positive constant $C=C(r)$.
@@
[19, Lemma 2.5] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines. Then
@@
2.4. Bergman kernel with sharp $\mathcal{B}$} conditions
In this subsection, we introduce a sharp $\mathcal{B}$} condition for Bergman kerenl, we firstly recall that the definition of $\mathcal{B}$}- systems.
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. For $z\in\overline{\Omega}$ near the boundary $\partial\Omega$, a family of functions $\mathcal{B}=\{b_{j}(z,\cdot)\}_{j=1}^{n}$ is called a $\mathcal{B}$-system at $z$ if there exist a neighbourhood $U$ of $z$ and a positive integer $m\geq 2$ such that for all $w\in U$,
$b_{1}(z,w):=\frac{1}{\delta(z,w)}\quad\text{and}\quad b_{j}(z,w):=\sum_{k=2}^{m}\left(\frac{A_{jk}(z)}{\delta(z,w)}\right)^{\frac{1}{k}},\quad\text{for }j=2,\ldots,n,$
@@ 2.4. Bergman kernel with sharp $\mathcal{B}$} conditions
$\delta(z,w):=|\delta_{\Omega}(z)|+|\delta_{\Omega}(w)|+|z_{1}-w_{1}|+\sum_{l=2}^{n}\sum_{s=2}^{m}A_{ls}(z)|z_{l}-w_{l}|^{s}$
@@
The Bergman kernel $K_{\Omega}$ is said to be of sharp $\mathcal{B}$-type at $z$ if $K_{\Omega}$ is of $\mathcal{B}$-type and has the sharp lower-bound on diagonal, i.e.
@@
We finish this section with a Forelli-Rudin estimate on domains with Bergman kernel satisfying the sharp $\mathcal{B}$} condition, which will be used to construct an atomic type decomposition.
@@
[19, Proposition 2.7] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain whose Bergman kernel $K_{\Omega}$ is of sharp $\mathcal{B}$-type. Assume that $\alpha\leq 0,\,\beta\geq 0,\,t+\alpha\geq 1$ and $1<\beta-2\alpha<2t-2$. Then the following inequality
@@ 3. Proof of main result
Let $r\in(0,1)$ be fixed, and $\{w_{k}\}$ be an $r$-lattice. For $0<p<\infty$ and $\{c_{k}\}\in l^{p}$, $\alpha+Np>1$. Set $f(z)=\sum_{k=1}^{\infty}c_{k}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}K_{\Omega}(z,w_{k})^{N}$,
$\|f\|_{A_{\alpha}^{p}}\lesssim 1.$
@@
Let $\Omega$ be a bounded pseudoconvex domain whose Bergman kernel $K_{\Omega}(z,w)$ satisfying sharp $\mathcal{B}$} condition. Assume that
@@
To prove the boundedness of $T_{q,s,t}$, we only prove the boundedness $S_{q,s,t}f(z)=K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t}f(w)dV(w)$. To do this, let $h(z)=K_{\Omega}(z,z)^{\sigma}$,
@@
Hence, lemma 3.1 that yields
@@
In order to prove (iv) $\Longleftrightarrow$ (v) in Theorem LABEL:restate, we need the following results.
@@
[19, Lemma 5.3] Let $1<p<\infty$ and $s\in\mathbb{R}$. Then the function $p_{\Omega}^{s}\widehat{\mu}_{r}\in L^{p}(\Omega)$ for some $r\in(0,1)$ if and only if $p_{\Omega}^{s}\widehat{\mu}_{R}\in L^{p}(\Omega)$ for all $R\in(0,1)$. Moreover, we have
@@
[19, Lemma 5.1] Suppose $r\in(0,1)$ and $0<p<\infty$. Then we have
@@
Thanks to the Montel’s theorem and Lemma 3.3, we need to prove (ii) $\implies$ (iii) $\implies$ (iv) $\implies$ (ii), (ii) $\implies$ (i).
@@
By Lemma 3.1, then $f\in A_{\alpha}^{p}$ and $\|f\|_{A_{\alpha}^{p}}^{p}\lesssim\sum_{k=0}^{\infty}|c_{k}|^{p}$. Since $J_{\mu}$ is bounded, hence
@@
(3.6) $\begin{aligned} \widehat{\mu}_{R}(z)&\approx p_{\Omega}(z)^{-2}\mu\left(B_{\Omega}(z,R)\right)\\ &\lesssim p_{\Omega}(a_{j})^{-2}\mu\left(B_{\Omega}(a_{j},r)\right)\approx\widehat{\mu}_{r}(a_{j})\end{aligned}$
Lemma 2.2, 2.4 and (3.6) yield that

### v2

# A note for Carleson measure on bounded $\mathbb{C}$}-convex domains

@@ Abstract.
Let $0<q<p<\infty$, $\Omega$ be a bounded $\mathbb{C}$}-convex domains in ${\mathbb{C}}^{n}$. We establish several equivalent characterizations for the boundedness of Carleson embedding $J_{\mu}:A_{\alpha}^{p}\hookrightarrow L^{q}(\mu)$ on $\Omega$ with sharp $\mathcal{B}$} type condition. Furthermore, we prove that the boundedness of $J_{\mu}$ is equivalent to its compactness.
@@ 1. Introduction and main result
The extension of $q$-Carleson measure to Bergman spaces has attracted considerable attention from many scholars, it was characterized by Hastings [18] and Oleinik [19] for $p\leq q$, Luecking [16] for $p>q$. In higher dimension, Cima-Wogen [7] and Luecking-Duren-Weir [14, 10] stuied the case of unit ball ${\mathbb{B}}^{n}$ by using the Euclidean geometry of ${\mathbb{B}}^{n}$ and conformal invariants (such as Bergman or Kobayashi balls) respectively. In order to extend the above results to other domains of ${\mathbb{C}}^{n}$, Cima-Mercer and Abate-Raissy-Saracco characterized the $q$-Carleson measure for $A^{p}$ on strongly pseudoconvex domains by using the Euclidean geometric quantities and intrinsic Kobayashi geometry respectively, see [1, 2, 3, 8] for the details. For the case of weakly pseudoconvex domains, Li-Liu-Wang [13] characterized the $2$-Carleson measurable for $A^{2}$ by using Kobayashi balls on bounded convex domains with smooth boundary of finite type. Bi-Su-Zhang [5] characterized the $2$-Carleson measure for $A^{2}$ and $A^{p}$ respectively on bounded $\mathbb{C}$}-convex domains.
Khanh-Liu-Thuc [11] introduced the sharp $\mathcal{B}$} type condition (see Sect.2.4 below) for Bergman kernel. Roughly speaking, Bergman kernel satisfies the sharp $\mathcal{B}$} type condition ensures that the Bergman projection is bounded on $L^{p}$ for $1<p<\infty$. More recently, there is a series of work involved function spaces and operator theory related such domains. For example, Khanh-Liu-Thuc-Tien [11, 12] characterized the boundedness and Schatten class of Bergman-Toeplitz operators for weighted spaces on weakly pseudoconvex domains with Bergman kernel satisfies sharp $\mathcal{B}$} type condition. Zhang [20] characterized the $q$-Carleson measure for the weighted Bergman spaces $A_{\alpha}^{p}$ on bounded $\mathbb{C}$}-convex domains with sharp $\mathcal{B}$} type condition, and what’s more, Zhang’s results extended the results in [5] to the case of $p\leq q$ and the weighted Bergman spaces $A_{\alpha}^{p}$, but for the case of $p>q$, Zhang didn’t characterize the full Carleson measure, Zhang’s result is stated as follows.
@@
[20, Theorem 1.8] Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel satisfies sharp $\mathcal{B}$} type condition, $\mu$ be a finite positive Borel measure on $\Omega$. Assume that $0<q<p<\infty$, if $\mu$ satisfies one of the following three conditions
@@
Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel of sharp $\mathcal{B}$} type. Let $\mu$ be a finite positive Borel measure, $0<q<p<\infty$. Then the following statements are equivalent.
@@ 2.2. Invariants
$\kappa_{\Omega}(x;v):=\inf\bigl\{|\xi|:f\in H(\mathbb{D},\Omega),\text{ with }f(0)=x,\;(df)_{0}(\xi)=v\bigr\},$
where $H({\mathbb{D}},\Omega)$ denotes the holomorphic section from unit disc $\mathbb{D}$} to $\Omega$. For any curve $\sigma:[a,b]\to\Omega$, the Kobayashi length is
@@
Suppose $\Omega$ is a bounded domain in ${\mathbb{C}}^{n}$, $z\in\Omega$ and $r\in(0,1)$. Let $B_{\Omega}(z,r)$ be the Kobayashi balls of centre $z$ and radius $\frac{1}{2}\log\frac{1+r}{1-r}$, i.e.
@@
For a bounded pseudoconvex domain $\Omega\subset{\mathbb{C}}^{n}$, Set $A^{2}$ be the Hilbert spaces consists of all $f\in{\mathcal{H}}(\Omega)$ such that $\|f\|_{A^{2}}^{2}=\int_{\Omega}|f(z)|^{2}dV(z)<\infty$, where $dV(z)$ is the ordinary Lebesgue measure. Riesz’s theorem yields that there exists a unique reproducing kernel function $K_{\Omega}(z,w)$ such that
@@ 2.3. Geometry of $\mathbb{C}$}-convex domains
and continuing this procedure we are led to an orthonormal basis
@@
[17, Theorem 1] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines, and $z\in\Omega$. Then for any $r\in(0,1)$,
@@
[5, Proposition 2.3] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. For any $r\in(0,1)$, we denote $R:=\frac{1}{2}(1+r)$. Then for any $z_{0}\in\Omega$ and $z\in B_{\Omega}(z_{0},r)$, we have
@@
The following cover type lemma plays a significant role in the proof of Theorem 1.2.
@@
[20, Lemma 2.4] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. Then for any $r\in(0,1)$, $z\in\Omega$ and $w\in B_{\Omega}(z,r)$, we have
@@
For $z\in\Omega$ and $r\in(0,1)$. Let $p_{\Omega}(z)=\prod\limits_{i=1}^{n}\eta_{i}(z)$, then by [20], for some positive constant $C=C(r)$ we have
(2.1) $\frac{1}{C}p_{\Omega}(z)\leq p_{\Omega}(w)\leq Cp_{\Omega}(z),\quad w\in B_{\Omega}(z,r).$
@@
[20, Lemma 2.5] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines. Then
@@
2.4. Bergman kernel with sharp $\mathcal{B}$} type conditions
In this subsection, we introduce a sharp $\mathcal{B}$} type condition for Bergman kernel, we firstly recall that the definition of $\mathcal{B}$}- systems.
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. For $z\in\overline{\Omega}$ near the boundary $\partial\Omega$, a family of functions $\mathcal{B}=\{b_{j}(z,\cdot)\}_{j=1}^{n}$ is called a $\mathcal{B}$-system at $z$ if there exist a neighborhood $U$ of $z$ and a positive integer $m\geq 2$ such that for all $w\in U$,
$b_{1}(z,w):=\frac{1}{\delta(z,w)}\quad\text{and}\quad b_{j}(z,w):=\sum_{k=2}^{m}\left(\frac{A_{jk}(z)}{\delta(z,w)}\right)^{\frac{1}{k}},\quad j=2,\ldots,n,$
@@ 2.4. Bergman kernel with sharp $\mathcal{B}$} type conditions
$\delta(z,w):=|\delta_{\Omega}(z)|+|\delta_{\Omega}(w)|+|z_{1}-w_{1}|+\sum_{l=2}^{n}\sum_{s=2}^{m}A_{ls}(z)|z_{l}-w_{l}|^{s}.$
@@
The Bergman kernel $K_{\Omega}$ is said to be of sharp $\mathcal{B}$ type at $z$ if $K_{\Omega}$ is of $\mathcal{B}$-type and has the sharp lower-bound on diagonal, i.e.
@@
We finish this section with a Forelli-Rudin estimate on domains with Bergman kernel satisfying the sharp $\mathcal{B}$} type condition, which will be used to construct an atomic type decomposition.
@@
[20, Proposition 2.7] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain whose Bergman kernel $K_{\Omega}$ is of sharp $\mathcal{B}$ type. Assume that $\alpha\leq 0,\,\beta\geq 0,\,t+\alpha\geq 1$ and $1<\beta-2\alpha<2t-2$. Then the following inequality
@@ 3. Proof of main result
Let $r\in(0,1)$ be fixed, and $\{w_{k}\}$ be an $r$-lattice. For $0<p<\infty$ and $\{c_{k}\}\in l^{p}$, $\alpha+Np>1$. Set $f(z)=\sum\limits_{k=1}^{\infty}c_{k}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}K_{\Omega}(z,w_{k})^{N}$, then we have
$\|f\|_{A_{\alpha}^{p}}\lesssim\|c_{k}\|_{l^{p}}\asymp 1.$
@@
Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded pseudoconvex domain whose Bergman kernel $K_{\Omega}(z,w)$ satisfies sharp $\mathcal{B}$} type condition. Assume that
@@
To prove the boundedness of $T_{q,s,t}$, we only prove the boundedness of $S_{q,s,t}f(z)=K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t}f(w)dV(w)$. To do this, let $h(z)=K_{\Omega}(z,z)^{\sigma}$,
@@
Hence, lemma 3.2 yields that
@@
In order to prove (iv) $\Longleftrightarrow$ (v) in Theorem 1.2, we need the following results.
@@
[20, Lemma 5.3] Let $1<p<\infty$ and $s\in\mathbb{R}$. Then the function $p_{\Omega}^{s}\widehat{\mu}_{r}\in L^{p}(\Omega)$ for some $r\in(0,1)$ if and only if $p_{\Omega}^{s}\widehat{\mu}_{R}\in L^{p}(\Omega)$ for all $R\in(0,1)$. Moreover, we have
@@
[20, Lemma 5.1] Suppose $r\in(0,1)$ and $0<p<\infty$. Then we have
@@
Thanks to the Montel’s theorem and Lemma 3.3, we only need to prove (ii) $\implies$ (iii) $\implies$ (iv) $\implies$ (ii), (ii) $\implies$ (i).
@@
By theorem 3.1, then $f\in A_{\alpha}^{p}$ and $\|f\|_{A_{\alpha}^{p}}^{p}\lesssim\sum_{k=0}^{\infty}|c_{k}|^{p}$. Since $J_{\mu}$ is bounded, hence
@@
(3.6) $\begin{aligned} \widehat{\mu}_{R}(z)&\approx p_{\Omega}(z)^{-2}\mu\left(B_{\Omega}(z,R)\right)\\ &\lesssim p_{\Omega}(a_{j})^{-2}\mu\left(B_{\Omega}(a_{j},r)\right)\approx\widehat{\mu}_{r}(a_{j}).\end{aligned}$
Lemma 2.2, lemma 2.4 and (3.6) yield that